OpenStudy (anonymous):

Find the range of each function for the domain {-4,-2,0,1.5,4}. 11. f(x)=-2x-3 12. f(x)=5x^2+4

OpenStudy (anonymous):

to find range put x=-4 and yor gonna get fx= -2*-4-3=8-3=5 n then put x=4 fx= -2*4-3=-8-3=-11 yor range is {5,11}

OpenStudy (anonymous):

find for the oder one, keep rocking ;)

OpenStudy (anonymous):

thanks thanks c:

OpenStudy (anonymous):

;)